OpenStudy (anonymous):

Hi i am trying to find the left and right and left Riemann sum on a ti 83. i have the program but my answer does not correlate with the calculator answer. does anyone know what the problem may be?

OpenStudy (nowhereman):

Which concrete problem are you calculating?

OpenStudy (anonymous):

im calculating for x^2+1 interval (0,3) with four rectangles for the left sum i got 8.906 but my calculator says its zero which is wrong

OpenStudy (anonymous):

hi nath,what Rieman sum you are working on?

OpenStudy (anonymous):

hey mark the question is above your response

OpenStudy (nowhereman):

I'm not quite familiar with that ti, what are you putting in there?

OpenStudy (anonymous):

for the y= i put in the function then i went to program selected integral and inputed the information but Im getting the wrong response

OpenStudy (anonymous):

ah ok, my pc keeps on hanging up thats why cant see them..lol

OpenStudy (mathteacher1729):

Here you go: http://www.math.psu.edu/dlittle/java/calculus/area.html Enter any function you like, find the left, right, and center sums. You can specify how many subintervals you'd like and it graphs it all. BEAUTIFUL applet. :)

OpenStudy (anonymous):

ty but i want yto do it on my own calculator so i can check my answer

OpenStudy (anonymous):

did you add you L+R?

OpenStudy (anonymous):

where should i add it

OpenStudy (anonymous):

on you calculation in the paper you calculated

OpenStudy (anonymous):

my calculation is right bc i checked my answer with the back of the book. there something wrong with my calculator

OpenStudy (anonymous):

approx A= (LHS+RHS)/2

OpenStudy (anonymous):

oh ok..what is the answer from the book?

OpenStudy (anonymous):

what is you LHS and the RHS? so that i can relate with my calculator also, im using a casio scintific calc

OpenStudy (anonymous):

mine is a graphing calcultor my left is 8.906 i didnt calculate my right

OpenStudy (anonymous):

calculate also your RHS

OpenStudy (anonymous):

my calculator directly computes for total Rieman sum, my calculator doesnt have a LEFT AND RIght sum

OpenStudy (anonymous):

ok nath i got the solution for you calculator now,,are you there

OpenStudy (anonymous):

use you calculator compute your delta x=(3-0)/4= .75. then on your calculator use the LHS Rieman sum= sigma (from 0 to 3)( (.75x)^2 +1) =11.875*.75=8.90625

OpenStudy (anonymous):

i will leave for you to solve for the RHS .......hope you get the idea nath